Find how many units of each mix are needed to meet daily requirement.
Daily req. should be = 30g of Protein, 16g of Fat, and 24g of Carb
Only 3 available mixes with the following grams of nutrient per unit(see below):


Mix X --- Protein 4 --- Fat 6 --- Carb 3
Mix Y --- Protein 6 --- Fat 1 --- Carb 2
Mix Z --- Protein 4 --- Fat 1 --- Carb 12

Answer by chessace(471)

L1: 4x+6y+4z=30 --> 2x+3y+2z=15
L2: 6x+y+z=16
L3: 3x+2y+12z=24
L4(L1*3): 6x+9y+6z=45
L5(L4-L2): 8y+5z=29
L6(2*L3 - L2): 3y+23z=32
L7(8*L6 - 3*L5): 169z=169
z=1
8y=24
y=3
6x+4=16
x=2